# Coresets from Trajectories: Selecting Data via Correlation of Loss Differences

### v1

@@ Abstract
Deep learning models achieve state-of-the-art performance across domains but face scalability challenges in real-time or resource-constrained scenarios. To address this, we propose Correlation of Loss Differences ($\mathtt{CLD}$), a simple and scalable metric for coreset selection that identifies the most impactful training samples by measuring their alignment with the loss trajectories of a held-out validation set. $\mathtt{CLD}$ is highly efficient, requiring only per-sample loss values computed at training checkpoints, and avoiding the costly gradient and curvature computations used in many existing subset selection methods. We develop a general theoretical framework that establishes convergence guarantees for $\mathtt{CLD}$-based coresets, demonstrating that the convergence error is upper-bounded by the alignment of the selected samples and the representativeness of the validation set. On CIFAR-100 and ImageNet-1k, $\mathtt{CLD}$-based coresets typically outperform or closely match state-of-the-art methods across subset sizes, and remain within 1% of more computationally expensive baselines even when not leading. $\mathtt{CLD}$ transfers effectively across architectures (ResNet, VGG, DenseNet), enabling proxy-to-target selection with $1\%$ degradation. Moreover, $\mathtt{CLD}$ is stable when using only early checkpoints, incurring negligible accuracy loss. Finally, $\mathtt{CLD}$ exhibits inherent bias reduction via per-class validation alignment, obviating the need for additional stratified sampling. Together, these properties make $\mathtt{CLD}$ a principled, efficient, stable, and transferable tool for scalable dataset optimization.
@@ 1 Introduction
Coresets support a range of applications including active learning (Coleman et al., 2020), neural architecture search (Na et al., 2021; Shim et al., 2021), dataset distillation (Cazenavette et al., 2022), and continual learning (Aljundi et al., 2019; Borsos et al., 2020). However, most existing approaches are either based on heuristic criteria unrelated to generalization (Toneva et al., 2018; Belouadah et al., 2020; Rebuffi et al., 2017), or expensive second-order or bilevel optimization (Killamsetty et al., 2021a; Pruthi et al., 2020; Garg & Roy, 2023; Killamsetty et al., 2021b; Xia et al., 2024; Borsos et al., 2020), limiting scalability.
@@
We validate these findings across CIFAR-100 and ImageNet-1k, where $\mathtt{CLD}$-selected coresets typically outperform or closely match state-of-the-art methods across a wide range of coreset sizes, and remain within 1% of more computationally expensive baselines even when not leading. Unlike these methods, $\mathtt{CLD}$ avoids costly gradients, pairwise similarities, or second-order statistics. It requires only per-sample loss values computed at training checkpoints, yielding significant gains in both computational and storage efficiency.
An additional strength of $\mathtt{CLD}$ is its robustness; the metric remains stable when computed using sparsely sampled training checkpoints and is consistent across random seeds, making it practical for large-scale or budgeted deployments. Furthermore, $\mathtt{CLD}$ coresets transfer effectively across architectures. Coresets computed using small proxy models (e.g., ResNet-18) generalize to larger models (e.g., ResNet-50, DenseNet) with performance drops consistently under $1\%$.
Figure 1: Correlation of Loss Differences (CLD) at a glance. Left: ImageNet-1k “Tiger” samples illustrating varying CLD scores. High-CLD samples (top row) closely track the validation loss trajectory, indicating informative and representative data. Low/negative-CLD samples (bottom row) significantly deviate, typically corresponding to atypical, ambiguous, or mislabeled examples. Right: Performance comparison of coresets formed by selecting equal-sized subsets of the highest 10% positive, lowest 10% negative, and 10% zero-valued CLD samples of ImageNet-1k on ResNet-18. Coresets with high-positive CLD samples achieve superior accuracy over various seeds.
@@ 2 Related Literature
Score-based methods select training samples according to predefined metrics, often in the feature space or based on model prediction confidence. Some methods score a sample by its distance to a class center (Rebuffi et al., 2017; Castro et al., 2018; Belouadah et al., 2020), to the class feature median (e.g., $\mathtt{Moderate}$ (Xia et al., 2022)), or to other samples (Sener & Savarese, 2018). $\mathtt{Cal}$ (Margatina et al., 2021) identifies contrastive examples via the KL divergence between predictive distributions, while $\mathtt{Herding}$ (Chen et al., 2010) selects representative samples using a kernel-based approach in Hilbert space. Other methods rely on prediction probabilities from a (possibly proxy) model. For example, $\mathtt{Forgetting}$ (Toneva et al., 2018) counts how often a sample is misclassified after previously being correct, $\mathtt{GraNd}$ and $\mathtt{EL2N}$ (Paul et al., 2021) rank examples by their early-training loss gradient norm or $l_{2}$ error norm, respectively, and $\mathtt{AUM}$ (Pleiss et al., 2020) scores the area under the margin across training to flag potential label issues. While these approaches are simple and avoid expensive gradient or Hessian computation, they often inherit biases from their heuristic scoring rules and offer no convergence guarantees. To mitigate bias, methods such as $\mathtt{CCS}$ (Zheng et al., 2022) employ stratified sampling for diversity, while $\mathbb{D}^{2}$-$\mathtt{Pruning}$ (Maharana et al., 2024) combines difficulty (prediction variance) and diversity (feature density) in a graph-based framework. Nonetheless, such strategies often make restrictive assumptions to avoid noisy samples (e.g., $\mathtt{CCS}$ discards up to $30\%$ of data) and still lack theoretical generalization guarantees.
@@
Optimization-based methods formulate coreset selection as an explicit optimization problem, often with provable convergence guarantees. $\mathtt{CRAIG}$ (Mirzasoleiman et al., 2020) and $\mathtt{GradMatch}$ (Killamsetty et al., 2021a) select samples that align with the full-data gradient direction, while $\mathtt{Glister}$ (Killamsetty et al., 2021b) maximizes held-out validation log-likelihood. Bilevel optimization (Borsos et al., 2020) has been used to leverage influence functions (Koh & Liang, 2017) for selecting samples with maximal generalization benefit, and $\mathtt{GraphCut}$ (Iyer et al., 2021) uses submodular information measures as the objective. Recently, $\mathtt{BoundarySet}$-$\mathtt{CCS}$ (Yang et al., 2024) minimized decision boundary reconstruction error while ensuring class diversity. While theoretically grounded, these approaches often require repeated optimization loops, making them computationally expensive for large datasets.
@@
Training property-based methods exploit the dynamics of training to assess sample importance. $\mathtt{SloCurv}$ (Garg & Roy, 2023) uses second-order loss curvature statistics to identify samples with better generalization potential, and $\mathtt{TracIn}$ (Pruthi et al., 2020) tracks gradient alignment with a validation set. $\mathtt{TDDS}$ (Zhang et al., 2024) extends this idea by projecting each sample’s gradient onto the accumulated gradient to quantify its true contribution, and by monitoring this projection across multiple iterations to account for fluctuations in importance over time. Although effective, such methods depend on costly first- or second-order statistics, limiting scalability.
@@
Scalable training property-based methods reduce this overhead by measuring per-sample dynamics using only forward-pass signals. $\mathtt{Dyn}$-$\mathtt{Unc}$ (He et al., 2024) summarizes the variability of the true-class probability over sliding windows and prunes by thresholding aggregated uncertainty after training. $\mathtt{DUAL}$ (Cho et al., 2025) combines this uncertainty score with a difficulty term and uses a pruning-ratio–adaptive Beta sampling schedule to reweight selection at high pruning ratios. In contrast, our method $\mathtt{CLD}$ ranks examples by how well their loss-difference trajectories align with the class-wise validation loss trajectory, providing an explicit generalization-alignment criterion. All three approaches admit identical minimal logging, one scalar per example per checkpoint (probability for $\mathtt{Dyn}$-$\mathtt{Unc}$$\mathtt{DUAL}$; loss for $\mathtt{CLD}$), so compute and storage are directly comparable. Unlike uncertainty-only methods, however, $\mathtt{CLD}$’s alignment objective supports a convergence guarantee (Section˜5), and it avoids ratio-specific sampling schedules and their hyperparameters; we report robustness at extreme cases (e.g., coresets of size $5\%$) under identical logging (Section˜6). In short, $\mathtt{CLD}$ combines the practicality of score-based metrics (low compute/storage via scalar logging) with a validation-aligned training-dynamics signal that admits a convergence guarantee.
@@ 4 Correlation of Loss Differences (CLD)
For every sample $\vec{z}$ we record the per-iteration change in loss during the model training run, and collect these $T$ increments in a loss-difference trajectory
@@
The validation-average trajectory is defined similarly as,
@@
where $\rho$ is the correlation metric. In our experiments, we employ Pearson correlation (Pearson, 1895) due to its scale invariance and computational simplicity. Intuitively, the $\mathtt{CLD}$ metric quantifies how well a training sample’s loss dynamics align with the aggregate loss trajectory of the validation set, which serves as a proxy for generalization behavior. Samples with higher $\mathtt{CLD}$ values are deemed more influential and can thus be prioritized for coreset construction. We investigate this hypothesis both theoretically and empirically in the following sections.
@@ 4.1 Coreset Selection Procedure
We first train a source model $\theta_{{\mathbf{S}}}$ on the full dataset $\mathbf{S}$}, recording per-epoch losses for all training and validation samples. We then compute $\vec{\Delta}_{m}$ and the class-specific average validation trajectory $\vec{\Delta}^{\prime}_{{\mathbf{V}},c}$ for each class $c$.
@@ 5 Theoretical Analysis of CLD-Coresets
These assumptions mirror those commonly adopted in analyses of training dynamics (Pruthi et al., 2020; Ilyas et al., 2022; Veen et al., 2020; Clemmensen & Kjærsgaard, 2022; Zhang et al., 2022); we merely state Assumption˜3 explicitly for transparency, even though it is typically invoked implicitly and is widely regarded as reasonable.
@@
Theorem˜1 shows that training on a high-$\mathtt{CLD}$ coreset achieves a similar bound, up to an additive deviation term $(B\sqrt{2\kappa}+\delta)^{2}$. This deviation captures the alignment of the coreset with the validation dynamics ($\kappa$), and the representativeness of the validation set ($\delta$).
The alignment term $\kappa$ reflects both the informativeness of selected samples and the size of the coreset. Higher $\mathtt{CLD}$ scores indicate stronger agreement with validation loss trajectories and thus tighter gradient alignment, reducing $\kappa$. Additionally, larger coresets more faithfully approximate full-data training dynamics, also lowering $\kappa$. When the coreset size is fixed, the theorem implies that selecting higher-$\mathtt{CLD}$ samples improves convergence by minimizing this deviation. Thus, $\mathtt{CLD}$-based selection emerges as a principled and necessary criterion for preserving the optimization behavior of full-data training.
@@ 6 Experimental Evaluation
We benchmark on CIFAR-100 (Krizhevsky et al., 2009) and ImageNet-1k (Russakovsky et al., 2015). CIFAR-100 has 50,00050{,}000 training and 10,00010{,}000 test images across $100$ classes; ImageNet-1k has $\sim\!1.28$M training images and a 50,00050{,}000-image validation set across 1,0001{,}000 classes. For each random seed, we form a classwise held-out validation split from the training data ($10\%$ for CIFAR-100; $1\%$ for ImageNet-1k), ensuring equal per-class representation; a different split is generated per seed, and the resulting train/validation partitions are reused across all baselines for fairness. Unless otherwise specified, ResNet-18 (He et al., 2016) is the default architecture for $\mathtt{CLD}$ scoring and for training on selected coresets. Coresets are constructed per seed in a class-balanced manner by selecting, within each class, the top-ranked samples under $\mathtt{CLD}$. Subset sizes range from $0.2\%$$100\%$ on CIFAR-100 and $\%1001$$0\%$ on ImageNet-1k. We report the mean and standard deviation over $5$ independent seeds.
Figure 2: Test accuracy (mean over five seeds) for representative coreset selection methods on CIFAR-100 and ImageNet-1k with ResNet-18. $\mathtt{CLD}$ consistently matches or outperforms baselines across dataset sizes. (X-axis uses a non-uniform coreset-size grid.) Color map: blues (score-based), oranges (optimization-based), greens (training-property-based), black ($\mathtt{CLD}$). Complete numerical results are available in Appendix˜C.
Baselines. We compare against representative state-of-the-art methods from three families: score-based ($\mathtt{Forgetting}$ (Toneva et al., 2018), $\mathtt{EL2N}$ (Paul et al., 2021), and $\mathtt{CCS}$ (Zheng et al., 2022) using $\mathtt{AUM}$ (Pleiss et al., 2020)), optimization-based ($\mathtt{Glister}$ (Killamsetty et al., 2021b), $\mathbb{D}^{2}$-$\mathtt{Pruning}$ (Maharana et al., 2024)), and training-property–based ($\mathtt{TDDS}$ (Zhang et al., 2024), $\mathtt{SloCurv}$ (Garg & Roy, 2023), $\mathtt{DUAL}$ (Cho et al., 2025)), plus $\mathtt{Random}$. We use implementations from the DeepCore (Guo et al., 2022) library when available, and otherwise rely on official GitHub repositories. All methods are run under a consistent training setup (40 pretraining epochs where required), without any additional fine-tuning or regularization. To ensure fairness, all baselines, including ours, select and train coresets using the same backbone (ResNet-18).

### v2

@@ Abstract
Deep learning models achieve state-of-the-art performance across domains but face scalability challenges in real-time or resource-constrained scenarios. To address this, we propose Correlation of Loss Differences ($\mathtt{CLD}$), a simple and scalable metric for coreset selection that identifies the most impactful training samples by measuring their alignment with the loss trajectories of a held-out validation set. $\mathtt{CLD}$ is highly efficient, requiring only per-sample loss values computed at training checkpoints, and avoiding the costly gradient and curvature computations used in many existing subset selection methods. We develop a general theoretical framework that establishes convergence guarantees for $\mathtt{CLD}$-based coresets, demonstrating that the convergence error is upper-bounded by the alignment of the selected samples and the representativeness of the validation set. On CIFAR-100 and ImageNet-1k, $\mathtt{CLD}$-based coresets typically outperform or closely match state-of-the-art methods across subset sizes, and remain within 1% of more computationally expensive baselines even when not leading. $\mathtt{CLD}$ transfers effectively across architectures (ResNet, VGG, DenseNet), enabling proxy-to-target selection with $1\%$ degradation. Moreover, $\mathtt{CLD}$ is stable when using only early checkpoints, incurring negligible accuracy loss. Finally, $\mathtt{CLD}$ exhibits inherent bias reduction via per-class validation alignment, obviating the need for additional stratified sampling. Together, these properties make $\mathtt{CLD}$ a principled, efficient, stable, and transferable tool for scalable dataset optimization.  [^1] The code is available on GitHub.
@@ 1 Introduction
Coresets support a range of applications including active learning (coreset_activelearning1), neural architecture search (coreset_nas1; coreset_nas2), dataset distillation (coreset_dc), and continual learning (coreset_cl_1; bilevelcoresets_borsos2020). However, most existing approaches are either based on heuristic criteria unrelated to generalization (forgetting_toneva2018; initial_beloudah2020; icarl_rebuffi2017), or expensive second-order or bilevel optimization (gradmatch_killamsetty2021; tracin_pruthi2020; slocurves_garg2023; glister_killamsetty2021; refinedcoreset_xia2024; bilevelcoresets_borsos2020), limiting scalability.
@@
We validate these findings across CIFAR-100 and ImageNet-1k, where $\mathtt{CLD}$-selected coresets typically outperform or closely match state-of-the-art methods across a wide range of coreset sizes, and remain within 1% of more computationally expensive baselines even when not leading. Unlike these methods, $\mathtt{CLD}$ requires only per-sample loss values, allowing it to avoid the costly gradients, Jacobians, or feature embeddings used by many influence- and similarity-based selectors. Concretely, the only computational overhead beyond a standard training run on the full dataset of size $N$ is a single forward pass at each checkpoint over a small, held-out validation set of size $Q$. Since the validation set is typically much smaller than the training set ($Q\ll N$), this lightweight approach yields significant gains in both computational and storage efficiency, which we quantify in our full cost analysis summarized in Table˜1.
An additional strength of $\mathtt{CLD}$ is its robustness; the metric remains stable when computed using sparsely sampled training checkpoints and is consistent across random seeds, making it practical for large-scale or budgeted deployments. Furthermore, $\mathtt{CLD}$ coresets transfer effectively across architectures. This is one of the key advantages of $\mathtt{CLD}$. Coresets computed using small proxy models (e.g., ResNet-18) generalize to larger models (e.g., ResNet-50, DenseNet) with performance drops consistently under $1\%$. Hence we can compute coresets using a smaller proxy model (e.g., ResNet-18) and reuse it to train larger target models (e.g., ResNet-50/VGG/DenseNet) with minimal loss in accuracy, while substantially reducing selection cost.
Figure 1: Correlation of Loss Differences (CLD) at a glance. Left: ImageNet‑1k “Tiger” samples illustrating varying CLD scores. High-CLD samples (top row) closely track the validation loss trajectory, indicating informative and representative data. Low/negative-CLD samples (bottom row) significantly deviate, typically corresponding to atypical, ambiguous, or mislabeled examples. Right: Performance comparison of coresets formed by selecting equal-sized subsets of the highest 10% positive, lowest 10% negative, and 10% zero-valued CLD samples of ImageNet-1k on ResNet-18. Coresets with high-positive CLD samples achieve superior accuracy over various seeds.
@@ 2 Related Literature
Score-based methods select training samples according to predefined metrics, often in the feature space or based on model prediction confidence. Some methods score a sample by its distance to a class center (icarl_rebuffi2017; endtoend_castro2018; initial_beloudah2020), to the class feature median (e.g., $\mathtt{Moderate}$ (xia2022moderate)), or to other samples (sener2018active). $\mathtt{Cal}$ (cal_margatina2021) identifies contrastive examples via the KL divergence between predictive distributions, while $\mathtt{Herding}$ (herding_chen2010) selects representative samples using a kernel-based approach in Hilbert space. Other methods rely on prediction probabilities from a (possibly proxy) model. For example, $\mathtt{Forgetting}$ (forgetting_toneva2018) counts how often a sample is misclassified after previously being correct, $\mathtt{GraNd}$ and $\mathtt{EL2N}$ (E2LN_grand_paul2021) rank examples by their early-training loss gradient norm or $l_{2}$ error norm, respectively, and $\mathtt{AUM}$ (pleiss2020_aum) scores the area under the margin across training to flag potential label issues. While these approaches are simple and avoid expensive gradient or Hessian computation, they often inherit biases from their heuristic scoring rules and offer no convergence guarantees. To mitigate bias, methods such as $\mathtt{CCS}$ (zheng_ccs) employ stratified sampling for diversity, while $\mathbb{D}^{2}$-$\mathtt{Pruning}$ (modelpred_d2_maharana2023) combines difficulty (prediction variance) and diversity (feature density) in a graph-based framework. Nonetheless, such strategies often make restrictive assumptions to avoid noisy samples (e.g., $\mathtt{CCS}$ discards up to $30\%$ of data) and still lack theoretical generalization guarantees.
@@
Optimization-based methods formulate coreset selection as an explicit optimization problem, often with provable convergence guarantees. $\mathtt{CRAIG}$ (craig_mirzasoleiman2020) and $\mathtt{GradMatch}$ (gradmatch_killamsetty2021) select samples that align with the full-data gradient direction, while $\mathtt{Glister}$ (glister_killamsetty2021) maximizes held-out validation log-likelihood. Bilevel optimization (bilevelcoresets_borsos2020) has been used to leverage influence functions (Koh2017) for selecting samples with maximal generalization benefit, and $\mathtt{GraphCut}$ (graphcut_iyer2021) uses submodular information measures as the objective. Recently, $\mathtt{BoundarySet}$-$\mathtt{CCS}$ (mindboundary_yang2024) minimized decision boundary reconstruction error while ensuring class diversity. While theoretically grounded, these approaches often require repeated optimization loops, making them computationally expensive for large datasets.
@@
Training property-based methods exploit the dynamics of training to assess sample importance. $\mathtt{SloCurv}$ (slocurves_garg2023) uses second-order loss curvature statistics to identify samples with better generalization potential, and $\mathtt{TracIn}$ (tracin_pruthi2020) tracks gradient alignment with a validation set. $\mathtt{TDDS}$ (zhang2024_tdds) extends this idea by projecting each sample’s gradient onto the accumulated gradient to quantify its true contribution, and by monitoring this projection across multiple iterations to account for fluctuations in importance over time. Although effective, such methods depend on costly first- or second-order statistics, limiting scalability.
@@
Scalable training property-based methods reduce this overhead by measuring per-sample dynamics using only forward-pass signals. $\mathtt{Dyn}$-$\mathtt{Unc}$ (uncertainity_he2024_dynunc) summarizes the variability of the true-class probability over sliding windows and prunes by thresholding aggregated uncertainty after training. $\mathtt{DUAL}$ (cho2025_DUAL) combines this uncertainty score with a difficulty term and uses a pruning-ratio–adaptive Beta sampling schedule to reweight selection at high pruning ratios. In contrast, our method $\mathtt{CLD}$ ranks examples by how well their loss-difference trajectories align with the class-wise validation loss trajectory, providing an explicit generalization-alignment criterion. All three approaches admit identical minimal logging, one scalar per example per checkpoint (probability for $\mathtt{Dyn}$-$\mathtt{Unc}$$\mathtt{DUAL}$; loss for $\mathtt{CLD}$), so compute and storage are directly comparable. Unlike uncertainty-only methods, however, $\mathtt{CLD}$’s alignment objective supports a convergence guarantee (Section˜5), and it avoids ratio-specific sampling schedules and their hyperparameters; we report robustness at extreme cases (e.g., coresets of size $5\%$) under identical logging (Section˜6). In short, $\mathtt{CLD}$ combines the practicality of score-based metrics (low compute/storage via scalar logging) with a validation-aligned training-dynamics signal that admits a convergence guarantee.
@@ 4 Correlation of Loss Differences (CLD)
For every sample $\vec{z}$ we record the per‑iteration change in loss during the model training run, and collect these $T$ increments in a loss‑difference trajectory
@@
The validation‑average trajectory is defined similarly as,
@@
where $\rho$ is the correlation metric. In our experiments, we employ Pearson correlation (pearsoncorr) due to its scale invariance and computational simplicity. Intuitively, the $\mathtt{CLD}$ metric quantifies how well a training sample’s loss dynamics align with the aggregate loss trajectory of the validation set, which serves as a proxy for generalization behavior. Samples with higher $\mathtt{CLD}$ values are deemed more influential and can thus be prioritized for coreset construction. We investigate this hypothesis both theoretically and empirically in the following sections.
@@ 4.1 Coreset Selection Procedure
We first train a source model $\theta_{{\mathbf{S}}}$ on the full dataset $\mathbf{S}$}, recording per-epoch losses for all training and validation samples. This logging piggybacks on the standard training loop; no additional passes over the training set are required. Consequently, all training samples are scored at the chosen checkpoints. We then compute $\vec{\Delta}_{m}$ and the class-specific average validation trajectory $\vec{\Delta}^{\prime}_{{\mathbf{V}},c}$ for each class $c$.
@@ 5 Theoretical Analysis of CLD-Coresets
These assumptions mirror those commonly adopted in analyses of training dynamics (tracin_pruthi2020; datamodels_ilyas2022; 2020_datarepresentativeness_validassump; 2022_datarepresentivity_validassump; 2022_deepactive_validassump); we merely state Assumption˜3 explicitly for transparency, even though it is typically invoked implicitly and is widely regarded as reasonable.
@@
Theorem˜1 shows that training on a high-$\mathtt{CLD}$ coreset achieves a similar bound, up to an additive deviation term $(B\sqrt{2\kappa}+\delta)^{2}$. This deviation captures the alignment of the coreset with the validation dynamics ($\kappa$), and the representativeness of the validation set ($\delta$), with $\delta=O(B/\sqrt{Q})$ decreasing in the validation size $Q$, and $\kappa$ decreasing as the CLD selection is tightened (smaller $\epsilon$) or as the coreset size $k$ increases (see Remark˜3 and Remark˜4 in Appendix˜B).
The alignment term $\kappa$ reflects both the informativeness of selected samples and the size of the coreset. Higher $\mathtt{CLD}$ scores indicate stronger agreement with validation loss trajectories and thus tighter gradient alignment, reducing $\kappa$. Additionally, larger coresets more faithfully approximate full-data training dynamics, also lowering $\kappa$.  We note that making $k$ extremely small can increase trajectory deviation $\|\delta_{t}\|$, which is reflected inside $\kappa$ (Remark 4). When the coreset size is fixed, the theorem implies that selecting higher-$\mathtt{CLD}$ samples improves convergence by minimizing this deviation. Thus, $\mathtt{CLD}$-based selection emerges as a principled and necessary criterion for preserving the optimization behavior of full-data training.
@@ 6 Experimental Evaluation
We benchmark on CIFAR-100 (cifar) and ImageNet-1k (imagenet). CIFAR-100 has 50,00050{,}000 training and 10,00010{,}000 test images across $100$ classes; ImageNet-1k has $\sim\!1.28$M training images and a 50,00050{,}000-image validation set across 1,0001{,}000 classes. For each random seed, we form a classwise held-out validation split from the training data ($10\%$ for CIFAR-100; $1\%$ for ImageNet-1k), ensuring equal per-class representation; a different split is generated per seed, and the resulting train/validation partitions are reused across all baselines for fairness. Unless otherwise specified, ResNet-18 (resnet) is the default architecture for $\mathtt{CLD}$ scoring and for training on selected coresets. Coresets are constructed per seed in a class-balanced manner by selecting, within each class, the top-ranked samples under $\mathtt{CLD}$. Subset sizes range from $0.2\%$$100\%$ on CIFAR-100 and $\%1001$$0\%$ on ImageNet-1k. We report the mean and standard deviation over $5$ independent seeds.
Figure 2: Test accuracy (mean over five seeds) for representative coreset selection methods on CIFAR-100 and ImageNet-1k with ResNet-18. $\mathtt{CLD}$ consistently matches or outperforms baselines across dataset sizes. (X-axis uses a non-uniform coreset-size grid.) Color map: blues (score-based), oranges (optimization-based), greens (training-property-based), black ($\mathtt{CLD}$). For reference, the full-data mean top-1 accuracy over five seeds is 70.95 on CIFAR-100 and 69.91 on ImageNet-1k. Complete numerical results are available in Appendix˜C.
Baselines. We compare against representative state-of-the-art methods from three families: score-based ($\mathtt{Forgetting}$ (forgetting_toneva2018), $\mathtt{EL2N}$ (E2LN_grand_paul2021), and $\mathtt{CCS}$ (zheng_ccs) using $\mathtt{AUM}$ (pleiss2020_aum)), optimization-based ($\mathtt{Glister}$ (glister_killamsetty2021), $\mathbb{D}^{2}$-$\mathtt{Pruning}$ (modelpred_d2_maharana2023)), and training-property–based ($\mathtt{TDDS}$ (zhang2024_tdds), $\mathtt{SloCurv}$ (slocurves_garg2023), $\mathtt{DUAL}$ (cho2025_DUAL)), plus $\mathtt{Random}$. We use implementations from the DeepCore (deepcore) library when available, and otherwise rely on official GitHub repositories. All methods are run under a consistent training setup (40 pretraining epochs where required), without any additional fine-tuning or regularization. To ensure fairness, all baselines, including ours, select and train coresets using the same backbone (ResNet-18).